OpenStudy (anonymous):

Identify the correct slope and y intercept of the equation 6x - 3y = 12.

OpenStudy (ivettef365):

just isolate the y, so the problem looks like this y = mx + b m = slope b = y-int

OpenStudy (anonymous):

lol idk, can you explain to me and give me the answer

OpenStudy (ivettef365):

yes, I could explain

OpenStudy (anonymous):

okay

OpenStudy (ivettef365):

6x -3y = 12 6x -6x -3y = -6x + 12 -3y = -6x + 12 --- --- --- -3 -3 -3

OpenStudy (ivettef365):

so you have y = 2x -4

OpenStudy (anonymous):

Given eq is 6x - 3y = 12. therefore 6x-12 = 3y or 3y = 6x-12 or y = 2x-4 comparinf it with y=mx+c we find slope m=2

OpenStudy (ivettef365):

now can you tell what is the slope and y -int

OpenStudy (anonymous):

whats the answer

OpenStudy (ivettef365):

can you tell what is the answer you have y = 2x - 4 compare to formula y = mx + b m = slope b = y-int

OpenStudy (anonymous):

idk what to do.

OpenStudy (ivettef365):

ok what is in the position of m ??

OpenStudy (ivettef365):

maybe you see it better this way y = (m) x (+ b) y = (2) x (- 4)

OpenStudy (anonymous):

good explaining ivette

OpenStudy (ivettef365):

thanks

OpenStudy (anonymous):

okay i got the answer.next question Write the equation of the line in slope-intercept form with a slope of 3 and a y-intercept of -2

OpenStudy (ivettef365):

ok same way you have the formula y = m x + b just plug m= slope = 3 and b = y-int = -2

OpenStudy (anonymous):

you can do this tevonn